Let x=2001^1002 - 2001^-1002 and y = 2001^1002 +2001^-1002. Find x^2 - y^2
Answer by stanbon(75887)

Let x=2001^1002 - 2001^-1002 and y = 2001^1002 +2001^-1002. Find x^2 - y^2
-------
Note:: x^2 - y^2 = (x+y)(x-y)
---
x+y = 2(2001^1002)
x-y = -2(2001^-1002)
------
(x+y)(x-y) = -4*1 = -4
-------------------
Cheers,
Stan H.
------------